write down standard equation for current and voltage in alternative current.
1
Expert's answer
2020-09-23T09:02:47-0400

"V(t)=V_p\sin(2\pi 
ft+\phi)" .............Voltage

"I(t)=I_p\sin(2\pi f t+\phi)" .............Current